OpenStudy (anonymous):

Find the critical point for ::::::::

OpenStudy (anonymous):

\[f(x)= x e ^{-x}\]

OpenStudy (anonymous):

the derivative is \[e ^{-x}-xe ^{-x}\] .. am i right ?

OpenStudy (anonymous):

Almost correct.

OpenStudy (anonymous):

what do u mean almost .. -_-?

OpenStudy (anonymous):

Actually it is correct, I didn't notice the xe^-x

OpenStudy (anonymous):

there is 1 critical point

OpenStudy (anonymous):

which is ... ?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

how did u know that .. from where u got it ..

OpenStudy (anonymous):

set it equal to 0 and solve

OpenStudy (anonymous):

okie .. can u show me ?

OpenStudy (anonymous):

e^−x−xe^−x=0

OpenStudy (anonymous):

okie i'l see what i can do with it o_o

OpenStudy (anonymous):

sankiyo anyway ^-^

OpenStudy (anonymous):

use ln ;)

OpenStudy (anonymous):

e^(-x)=xe^(-x)